camo15:

The polynomial below is a perfect square trinomial of the form A^2-2AB+B^2. 4x^2-16x+16. True or False

Mercury:

well, let's start by assuming it *is* a perfect square trinomial if A^2 - 2AB + B^2 = 4x^2 - 16x + 16, what do A and B have to be? as a hint, if it's a perfect square trinomial, then: A^2 = 4x^2 B^2 = 16 -2AB = -16x are there any such A and B values that will make all three equations true?